The amount that a sample of eight customers spent for lunch ($) at a fast-food restaurant is shown below:

67, 52, 86, 52, 88, 32, 28, 43

a. Find the sample mean ???

b. Find the sample standard deviation S

c. Construct a 95% confidence interval estimate for the population mean ?
Answer #1

Solution:

Given that

a )The mean of sample is   \bar x

\sumx/n =  (67+ 52+ 86+ 52+ 88+ 32+ 28+ 43 / 8 )

= 448 / 8

= 56

The sample mean is 56

b ) Sample standard deviation is s

s = \sqrt1/(n-1)\sum(x - \bar x )2

=  \sqrt1/7 (67 - 56 )2+ (52 - 56 )2+ (86 - 56 )2+ (52- 56 )2 +( 88 - 56 )2+ (32 - 56 )2 + (28 - 56 )2+ (43- 56 )2)

= \sqrt1/7 (121+16+900+16+1024+576+784+169)

= \sqrt3606/7

= \sqrt515.14

= 22.70

Sample standard deviation = 22.70

c ) \bar x = 56

s = 22.70

n = 8

Degrees of freedom = df = n - 1 = 7

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,7 = 2.365

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.365 * (22.70 / \sqrt8) = 18.9

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

56 - 18.98 < \mu < 56 + 18.98

37.02 < \mu < 74.98

(37.02, 74.98 )